6 3/4 x 5 1/3=
dont know how to solve?

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

6 3/4 x 5 1/3
----
= (27/4)*(16/3)
----
Cancel any factor common to a numerator and a denominator:
---
= (9/1)(4/1)
----
= 36
======
Cheers,
Stan H.

Answer by rfer(16322)

change to improper fractions
multiply
then reduce
27/4*16/3=432/12=36